OpenStudy (anonymous):

cscU -1/cscU+1= 1-sinU/1+sinU

OpenStudy (anonymous):

Trying to find the identities

OpenStudy (anonymous):

\(\large\cal\color{red}{Welcome}~\color{red}{to}~\color{red}{Openstudy!}\)

OpenStudy (anonymous):

Since csc(u) = 1/sin(u) and cot(u) = cos(u)/sin(u), we have... LHS = (1/sin(u)) - cos(u)/(sin(u)) = (1 - cos(u))/sin(u) Then, multiply the top and bottom by 1 + cos(u) to get... (1 - cos(u))/sin(u) * (1 + cos(u))/(1 + cos(u)) = (1 - cos²(u))/(sin(u)(1 + cos(u))) = (1 - cos²(u))/(sin(u)(1 + cos(u))) Finally, by noting that 1 - cos²(u) = sin²(u), we can simplify the expression to... sin²(u)/(sin(u)(1 + cos(u))) = sin(u)/(1 + cos(u)) = RHS Hope this helped! :D